# A Practitioner’s Guide to Kolmogorov–Arnold Networks

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have recently emerged as a promising alternative to traditional Multilayer Perceptrons (MLPs), inspired by the Kolmogorov-Arnold representation theorem. Unlike MLPs, which use fixed activation functions on nodes, KANs employ learnable univariate basis functions on edges, offering enhanced expressivity and interpretability. This review provides a systematic and comprehensive overview of the rapidly expanding KAN landscape, moving beyond simple performance comparisons to offer a structured synthesis of theoretical foundations, architectural variants, and practical implementation strategies. By collecting and categorizing a vast array of open-source implementations, we map the vibrant ecosystem supporting KAN development. We begin by bridging the conceptual gap between KANs and MLPs, establishing their formal equivalence and highlighting the superior parameter efficiency of the KAN formulation. A central theme of our review is the critical role of the basis function; we survey a wide array of choices—including B-splines, Chebyshev and Jacobi polynomials, ReLU compositions, Gaussian RBFs, and Fourier series—and analyze their respective trade-offs in terms of smoothness, locality, and computational cost. We then categorize recent advancements into a clear roadmap, covering techniques for improving accuracy, efficiency, and regularization. Key topics include physics-informed loss design, adaptive sampling, domain decomposition, hybrid architectures, and specialized methods for handling discontinuities. Finally, we provide a practical “Choose-Your-KAN” guide to help practitioners select appropriate architectures, and we conclude by identifying current research gaps. We argue for a shift away from simplistic KAN-vs-MLP benchmarks toward a more methodical, basis-centric exploration of this rich architectural paradigm. The associated GitHub repository (https://github.com/AmirNoori68/kan-review) complements this paper and serves as a structured reference for ongoing KAN research.
@@ 1 Introduction
Kolmogorov–Arnold Networks [27] introduce a distinct parameterization aimed at overcoming several limitations of MLPs. Instead of relying on fixed nonlinearities, KANs place learnable univariate functions on edges, inspired by the Kolmogorov representation theorem, which decomposes multivariate mappings into sums of one-dimensional transforms. The original implementation employs B-spline bases [27], but the framework is inherently modular: Chebyshev polynomials, Gaussian kernels, and other bases can be substituted to suit the problem structure. We regard this flexibility as an asset rather than a complication, since the choice of basis family directly governs smoothness, locality, and spectral behavior—thereby shaping both expressivity and interpretability. Consequently, the combination of adaptive basis learning with the freedom to select different basis families represent a major advantage of the KAN approach. Beyond basis design, KANs encompass a broad landscape of architectural extensions, optimization strategies, and theoretical developments—an overview of which is outlined in Table 1 to guide the reader through the remainder of this paper.
@@
1 Introduction 2 MLP Basics 3 KAT and KAN 4 Bridging KANs and MLPs 5 How KANs Extend the MLPs Landscape 6 Basis Functions $\triangleright$ 6.1 B-spline $\triangleright$ 6.2 Chebyshev Polynomial $\triangleright$ 6.3 ReLU $\triangleright$ 6.4 Jacobi Polynomials $\triangleright$ 6.5 Gaussian RBF $\triangleright$ 6.6 Fourier $\triangleright$ 6.7 Wavelet $\triangleright$ 6.8 Finite-Basis $\triangleright$ 6.9 SincKAN 7 Accuracy Improvement $\triangleright$ 7.1 Physics Constraints & Loss Design $\triangleright$ 7.2 Adaptive Sampling & Grids $\triangleright$ 7.3 Domain Decomposition $\triangleright$ 7.4 Function Decomposition $\triangleright$ 7.5 Hybrid/Ensemble & Data $\triangleright$ 7.6 Sequence/Attention Hybrids $\triangleright$ 7.7 Discontinuities & Sharp Gradients $\triangleright$ 7.8 Optimization & Adaptive Training 8 Efficiency Improvement $\triangleright$ 8.1 Parallelism, GPU, and JAX Engineering $\triangleright$ 8.2 Matrix Optimization & Efficient Bases 9 Sparsity & Regularization 10 Scaling Laws & Convergence $\triangleright$ 10.1 Theoretical Approximation Rates $\triangleright$ 10.2 Spectral Bias $\triangleright$ 10.3 NTK-Based Convergence $\triangleright$ 10.4 Practical Trade-offs 11 Practical “Choose–Your–KAN” Guide 12 Current Gaps and Path Forward 13 Conclusion
@@
A number of recent surveys have appeared on KANs, each highlighting different facets of this emerging architecture. Notable contributions include those by Andrade et al. [67], Basina et al. [68], Beatrize et al. [69], Faroughi et al. [70], Kilani et al. [71], Hou et al. [72], Dutta et al. [73], Essahraui et al. [74], and Somvanshi et al. [75], which together provide valuable entry points into the literature. Our review builds on these efforts by seeking to offer a more systematic and comprehensive perspective: rather than cataloguing studies in isolation, we integrate theoretical, architectural, optimization, and application viewpoints into a structured roadmap. By combining comparative analysis, methodological insights, and practical guidance (Sections 611), our goal is to provide a resource that complements existing surveys and serves readers aiming to both understand and apply KANs.
In parallel, KANs are rapidly expanding on open-access platforms, with diverse implementations and variants shared by the community. Curated repositories such as [76] [^3] https://github.com/mintisan/awesome-kan track the latest developments across different fields, making this growth more accessible to readers. Table 3 summarizes representative GitHub repositories that illustrate this fast-growing ecosystem in regression, function approximation, and PDE solving. These resources are also referenced in footnotes throughout this study when further details of specific methods are discussed.
@@
Repository (https://github.com/…) Description Ref.
KindXiaoming/pykan Official PyKAN for “KAN” and “KAN 2.0”. [77]
afrah/pinn_learnable_activation Compares various KAN bases vs. MLP on PDEs [31]
1ssb/torchkan Simplified PyTorch KAN with variants [78]
mintisan/awesome-kan Curated list of KAN resources, projects, and papers. [76]
sidhu2690/Deep-KAN Spline-KAN examples and a PyPI package. [79]
sidhu2690/RBF-KAN RBF-KAN examples [80]
yu-rp/KANbeFair Fair benchmarking of KANs vs MLPs. [28]
Blealtan/efficient-kan Efficient PyTorch implementation of KAN. [81]
srigas/jaxKAN JAX-based KAN package with grid extension support. [82]
ZiyaoLi/fast-kan FastKAN using RBFs. [83]
AthanasiosDelis/faster-kan Using SWitch Activation Function [84]
Indoxer/LKAN Implementations of KAN variations. [85]
pnnl/neuromancer (fbkans branch) Parametric constrained optimization. [86]
quiqi/relu_kan Minimal ReLU-KAN. [87]
OSU-STARLAB/MatrixKAN Matrix-parallelized KAN. [88]
Iri-sated/PowerMLP MLP-type network with KAN-level expressiveness. [40]
GistNoesis/FourierKAN Fourier-based KAN layer. [89]
GistNoesis/FusedFourierKAN Optimized FourierKAN with fused GPU kernels [90]
alirezaafzalaghaei/fKAN Fractional KAN using Jacobi functions. [91]
alirezaafzalaghaei/rKAN Rational KAN (Padé/Jacobi rational designs). [92]
M-Wolff/CVKAN Complex-valued KANs. [93]
DUCH714/SincKAN Sinc-based KAN with PINN applications. [41]
SynodicMonth/ChebyKAN Chebyshev polynomial-based KAN variant. [94]
Boris-73-TA/OrthogPolyKANs Orthogonal polynomial-based KAN implementations. [95]
kolmogorovArnoldFourierNetwork/kaf_act PyTorch activation combining with RFF. [96]
kolmogorovArnoldFourierNetwork/KAF Kolmogorov-Arnold Fourier Networks. [63]
kelvinhkcs/HRKAN Higher-order ReLU-KANs. [97]
yizheng-wang/KINN PIKAN for solid mechanics PDEs. [61]
Ali-Stanford/KAN_PointNet_CFD Jacobi-based network for CFD predictions. [62]
Jinfeng-Xu/FKAN-GCF FourierKAN-GCF for graph filtering. [98]
@@
jdtoscano94/KKANs_PIML Kurkova-KANs combining MLP with basis functions. [44]
Zhangyanbo/MLP-KAN MLP-augmented KAN activations. [99]
Adamdad/kat Kolmogorov-Arnold Transformer. [100]
YihongDong/FAN Fourier Analysis Network (FAN). [101]
seydi1370/Basis_Functions Polynomial bases for KANs (comparative study). [102]
zavareh1/Wav-KAN Wav-KAN: wavelet-based KANs. [103]
Jim137/qkan Quantum variational activations and pruning tools. [43]
liouvill/KAN-Converge Additive & hybrid KANs for convergence-rate experiments [104]
hoangthangta/BSRBF_KAN Combines B-splines (BS) and radial basis functions (RBF). [32]
wmdataphys/Bayesian-HR-KAN Introduces Bayesian higher-order ReLU-KANs. [105]
zhang-zhuo001/Legend-KINN Legendre polynomial–based KAN. [66]
DiabAbu/DeepOKAN Deep Operator Network based on KAN. [46]
DENG-MIT/LeanKAN A memory-efficient Kolmogorov–Arnold Network. [106]
@@
Organization (Table 1). Section 2 reviews MLP basics. Section 3 introduces the Kolmogorov–Arnold theorem (KAT) and formalizes KAN layers. Section 4 connects KANs to classical neural networks. Section 6 surveys basis families and computational trade-offs. Section 7 covers accuracy-improvement strategies, and Section 8 addresses efficiency via parallelism/GPU/JAX and matrix optimization. Section 9 treats sparsity, regularization, and Bayesian variants, while Section 10 discusses scaling laws and convergence. Section 11 presents the practical “Choose–Your–KAN” guide. We conclude with final remarks in Section 13.

### v2

@@ Abstract
The so-called Kolmogorov–Arnold Networks (KANs), whose design is merely inspired, rather than dictated, by the Kolmogorov superposition theorem, have emerged as a promising alternative to traditional Multilayer Perceptrons (MLPs). This review provides a systematic and comprehensive overview of the rapidly expanding KAN landscape. By collecting and categorizing a large set of open-source implementations, we map the vibrant ecosystem supporting modern KAN development. We organize the review around four core themes:
(i) presenting a precise history of Kolmogorov’s superposition theory toward neural-network formulations (Sec. 3);
(ii) establishing the formal equivalence between KANs and MLPs (Secs. 4,5);
(iii) analyzing the critical role of basis functions (Sec. 6); and
(iv) organizing recent advancements in accuracy, efficiency, regularization, and convergence (Secs. 7, 8, 9, 10).
Finally, we provide a practical Choose-Your-KAN guide (Sec. 11) to assist practitioners in selecting appropriate architectures, and we close by identifying current research gaps and future directions (Sec. 12). The associated GitHub repository (https://github.com/AmirNoori68/kan-review) complements this paper and serves as a structured reference for ongoing KAN research.
@@ 1 Introduction
KANs [26] replace fixed activations with learnable univariate functions on edges, motivated by the Kolmogorov representation theorem. The original model uses B-spline bases [26], but the framework is modular: Chebyshev polynomials, Gaussian kernels, and other basis families can be substituted as needed. This flexibility is central, since the chosen basis controls smoothness, locality, and spectral behavior, directly affecting expressivity and interpretability. The combination of adaptive basis learning and interchangeable basis families is therefore a core advantage of KANs. KANs also include a growing set of architectural variants, optimization strategies, and theoretical developments.
A number of recent surveys have appeared on KANs, each highlighting different facets of this emerging architecture. Notable contributions include those by Andrade et al. [27], Basina et al. [28], Beatrize et al. [29], Faroughi et al. [30], Kilani et al. [31], Hou et al. [32], Dutta et al. [33], Essahraui et al. [34], and Somvanshi et al. [35], which together provide valuable entry points into the literature. Our review builds on these efforts by seeking to offer a more systematic and comprehensive perspective: rather than cataloguing studies in isolation, we integrate theoretical, architectural, optimization, and application viewpoints into a structured roadmap. By combining comparative analysis, methodological insights, and practical guidance, our goal is to provide a resource that complements existing surveys and serves readers aiming to both understand and apply KANs. Table 1 presents a structured roadmap of the paper’s sections and themes.
@@
1 Introduction 2 MLP Basics 3 KST Toward Neural Networks 4 Bridging KANs and MLPs 5 How KANs Extend the MLPs Landscape 6 Basis Functions $\triangleright$ 6.1 B-spline $\triangleright$ 6.2 Chebyshev Polynomial $\triangleright$ 6.3 ReLU $\triangleright$ 6.4 Jacobi and General Polynomials $\triangleright$ 6.5 Gaussian RBF $\triangleright$ 6.6 Fourier $\triangleright$ 6.7 Wavelet $\triangleright$ 6.8 Finite-Basis $\triangleright$ 6.9 SincKAN 7 Accuracy Improvement $\triangleright$ 7.1 Physics Constraints & Loss Design $\triangleright$ 7.2 Adaptive Sampling & Grids $\triangleright$ 7.3 Domain Decomposition $\triangleright$ 7.4 Function Decomposition $\triangleright$ 7.5 Hybrid/Ensemble & Data $\triangleright$ 7.6 Sequence/Attention Hybrids $\triangleright$ 7.7 Discontinuities & Sharp Gradients $\triangleright$ 7.8 Optimization & Adaptive Training 8 Efficiency Improvement $\triangleright$ 8.1 Parallelism, GPU, and JAX Engineering $\triangleright$ 8.2 Matrix Optimization & Efficient Bases 9 Sparsity & Regularization 10 Scaling Laws & Convergence $\triangleright$ 10.1 Theoretical Approximation Rates $\triangleright$ 10.2 Spectral Bias $\triangleright$ 10.3 NTK-Based Convergence $\triangleright$ 10.4 Practical Trade-offs 11 Practical “Choose–Your–KAN” Guide 12 Current Gaps and Path Forward 13 Conclusion
@@
Repository Description Ref.
…/pykan Official PyKAN for “KAN” and “KAN 2.0”. [113]
…/Anant-Net High-dimensional PDE solver with tensor sweeps. [256]
…/RGA-KANs Deep cPIKANs with variance-preserving initialization. [262]
…/pinn_learnable_activation Compares various KAN bases vs. MLP on PDEs [39]
…/torchkan Simplified PyTorch KAN with variants [114]
…/awesome-kan Curated list of KAN resources, projects, and papers. [75]
…/Deep-KAN Spline-KAN examples and a PyPI package. [115]
…/RBF-KAN RBF-KAN examples [116]
…/KANbeFair Fair benchmarking of KANs vs MLPs. [36]
…/efficient-kan Efficient PyTorch implementation of KAN. [117]
…/spikans Separable PIKAN (SPIKAN). [190]
…/jaxKAN JAX-based KAN package with grid extension support. [259]
…/fast-kan FastKAN using RBFs. [119]
…/faster-kan Using SWitch Activation Function [120]
…/LKAN Implementations of KAN variations. [121]
…/neuromancer>->fbkans Parametric constrained optimization. [122]
…/relu_kan Minimal ReLU-KAN. [123]
…/MatrixKAN Matrix-parallelized KAN. [124]
…/PowerMLP MLP-type network with KAN-level expressiveness. [48]
…/FourierKAN Fourier-based KAN layer. [125]
…/FusedFourierKAN Optimized FourierKAN with fused GPU kernels [126]
…/fKAN Fractional KAN using Jacobi functions. [127]
…/rKAN Rational KAN (Padé/Jacobi rational designs). [128]
…/CVKAN Complex-valued KANs. [129]
…/SincKAN Sinc-based KAN with PINN applications. [49]
…/ChebyKAN Chebyshev polynomial-based KAN variant. [130]
…/OrthogPolyKANs Orthogonal polynomial-based KAN implementations. [131]
…/kaf_act PyTorch activation combining with RFF. [132]
…/KAF Kolmogorov-Arnold Fourier Networks. [71]
…/HRKAN Higher-order ReLU-KANs. [133]
…/KINN PIKAN for solid mechanics PDEs. [69]
…/KAN_PointNet_CFD Jacobi-based network for CFD predictions. [70]
…/FKAN-GCF FourierKAN-GCF for graph filtering. [134]
@@
…/KKANs_PIML Kurkova-KANs combining MLP with basis functions. [52]
…/MLP-KAN MLP-augmented KAN activations. [135]
…/kat Kolmogorov-Arnold Transformer. [136]
…/FAN Fourier Analysis Network (FAN). [137]
…/Basis_Functions Polynomial bases for KANs (comparative study). [138]
…/Wav-KAN Wav-KAN: wavelet-based KANs. [139]
…/qkan Quantum variational activations and pruning tools. [51]
…/KAN-Converge Additive & hybrid KANs for convergence-rate experiments [140]
…/BSRBF_KAN Combines B-splines and radial basis functions. [40]
…/Bayesian-HR-KAN Introduces Bayesian higher-order ReLU-KANs. [141]
…/Legend-KINN Legendre polynomial–based KAN. [74]
…/DeepOKAN Deep Operator Network based on KAN. [54]
…/LeanKAN A memory-efficient Kolmogorov–Arnold Network. [142]
…/KANQAS_code KANQAS: KAN for Quantum Architecture Search. [257]
…/pkan Probabilistic KAN via divisive data re-sorting. [258]
…/pihnn Holomorphic KAN/PIHNN framework for solving PDEs. [255]
…/lmkan Lookup-based KAN for fast high-dim mappings. [268]
…/KAN_Initialization_Schemes Initialization schemes for spline-based KANs. [276]
…/mlp-kan KAN vs. MLP for PDEs in DeepONet/GNS frameworks. [267]
KANs are rapidly expanding on open-access platforms, with diverse implementations and variants shared by the community. Curated repositories such as [75] [^3] https://github.com/mintisan/awesome-kan track the latest developments across different fields, making this growth more accessible to readers. Table LABEL:githubs summarizes representative GitHub repositories that illustrate this fast-growing ecosystem in regression, function approximation, and PDE solving. These resources are also referenced in footnotes throughout this study when further details of specific methods are discussed.
@@
Organization (Table 1). Section 2 reviews MLP basics. Section 3 introduces Kolmogorov Superposition Theory and its refinements. Section 4 connects KANs to classical neural networks. Section 6 surveys basis families and computational trade-offs. Section 7 covers accuracy-improvement strategies, and Section 8 addresses efficiency via parallelism/GPU/JAX and matrix optimization. Section 9 treats sparsity, regularization, and Bayesian variants, while Section 10 discusses scaling laws and convergence. Section 11 presents the practical “Choose–Your–KAN” guide. We conclude with final remarks in Section 13.